OpenStudy (anonymous):

Can I get some more help? I'm confused. √2x-7=4

jagr2713 (jagr2713):

\[\huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\] @Taylorcheer98 what do you think we do first

OpenStudy (anonymous):

The 2x-7?

sammixboo (sammixboo):

Wait, is it \(\sqrt{2x-7} = 4\) or \(\sqrt{2x} - 7 = 4\) ?

jagr2713 (jagr2713):

^^

OpenStudy (anonymous):

It's the first one. What is the difference between the two?

TheSmartOne (thesmartone):

A huge difference.

TheSmartOne (thesmartone):

It could get you a wrong answer. Always good to clarify :)

OpenStudy (anonymous):

Could you tell me the difference between the two

jagr2713 (jagr2713):

first we square both sides of the equation.

sammixboo (sammixboo):

Well, we know that \(\sqrt{2x-7}=4\) is the same as \(2x-7 = 4^2\) , right?

jagr2713 (jagr2713):

you got this @sammixboo ?

sammixboo (sammixboo):

If you want to continue I'll leave :o

sammixboo (sammixboo):

=3

jagr2713 (jagr2713):

its ok

OpenStudy (anonymous):

Yes

TheSmartOne (thesmartone):

Well @Taylorcheer98 the first one is easier to solve, the second takes more steps. And you will get 2 different answers. Because notice how the 7 or could not be under the squareroot?

OpenStudy (anonymous):

light.

OpenStudy (anonymous):

what he said

OpenStudy (anonymous):

Yes

sammixboo (sammixboo):

OK so we have \(2x-7 = 4^2\) Let's do the exponent, \(4^2\) first. Do you know what that is equal to? \(4^2\rightarrow 4\times 4\)

OpenStudy (anonymous):

16?

sammixboo (sammixboo):

Right so \(2x - 7 = 16\) Do you know what to do next? Solve this like a regular equation

OpenStudy (anonymous):

Ok, let me solve it

sammixboo (sammixboo):

OK tell us what you get, so we can check your answer x3

OpenStudy (anonymous):

23/2?

sammixboo (sammixboo):

Correct!!

OpenStudy (anonymous):

Thank you I understand it better'!

sammixboo (sammixboo):

=3 if you need anymore help tag me @sammixboo

OpenStudy (anonymous):

I definitely will.

sammixboo (sammixboo):

Okie dokie! Welcome to OpenStudy, once again! If you have any concerns on how to use the site, Private Message me! I will reply ASAP

OpenStudy (anonymous):

Thank you so much! I appreciate that.

TheSmartOne (thesmartone):

:)